You’ve observed the following returns on Crash-n-Burn Computer’s stock over the past five years: 5 percent, -15 percent, 26 percent, 21 percent, and 16 percent.

  

Requirement 1:
What was the arithmetic average return on Crash-n-Burn’s stock over this five year period?
(Click to select)  12.61%  11.55%  10.60%  13.25%  8.59%

   

Requirement 2:
(a)

What was the variance of Crash-n-Burn’s returns over this period? (Do not round intermediate calculations.)

(Click to select)  0.02414  0.02149  0.02122  0.02741  0.02653

   

(b)

What was the standard deviation of Crash-n-Burn’s returns over this period? (Do not round intermediate calculations.)

(Click to select)  16.56%  14.57%  13.19%  16.29%  19.38%
Answer #1

1) To find the average return, we sum all the returns and divide by the number of returns, so:

Average return = (0.05 – 0.15 + 0.26 + 0.21 + 0.16)/5 = 0.1060, or 10.60%

2) (a) Using the equation to calculate variance, we find:

Variance = 1/4[(0.05 – 0.1060)2 + (– 0.15 – 0.1060)2 + ( 0.26 – 0.1060)2 + ( 0.21 – 0.1060)2 + ( 0.16 – 0.1060)2]

Variance = 0.02653

(b) Standard deviation = (0.02653)1/2 = 0.1629, or 16.29%